OpenStudy (anonymous):

Part B: Determine three different solutions for this equation. 7x + y = 5 But the equation turns over to be y=-7x+5. You solve the equation how?

OpenStudy (triciaal):

is this all of it? when x = 0, y= 5 when y = 0, x = 5/7 maximum value when x = -7

OpenStudy (anonymous):

No there is more but I already answered the Part A. I also need help with Part C. If you can help?

OpenStudy (anonymous):

An equation is shown below: 7x + y = 5 Part A: Explain how you will show all of the solutions that satisfy this equation. Part B: Determine three different solutions for this equation. Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

you still there?

OpenStudy (triciaal):

yes no clue what is an inconsistent equation?

OpenStudy (anonymous):

lol An inconsistent equation is an equation that has no solution.

OpenStudy (triciaal):

thanks

OpenStudy (anonymous):

no problem :) So you dont know how to answer this question either?

OpenStudy (triciaal):

no

OpenStudy (anonymous):

Ok. Oh well! Thanks anyways!